# Match, Compare, or Select? An Investigation of Large Language Models for Entity Matching

### v1

@@ 6 Conclusion
In this paper, we investigate three strategies for entity matching with LLMs to bridge the gap between local matching and global consistency of ER. Our research shows that incorporating record interactions is essential for LLM-based entity matching. By examining the effect of broad LLMs on these strategies, we further design a ComEM framework that integrates the advantages of multiple strategies and LLMs. The effectiveness and cost efficiency of ComEM highlight the importance of task decomposition and LLM composition, opening up new avenues for entity matching with LLMs.
@@
This study aims to investigate different strategies for entity matching with LLMs. We conducted experiments using 1 commercial LLM and 9 open-source LLMs to provide a broad base for our analyses. Nevertheless, the dynamic update of LLMs and the influence of more powerful LLMs need to be further investigated. Besides, prompt engineering or fine-tuned approaches for different strategies are also a worthwhile direction. We have also demonstrated the effectiveness of the composition framework in entity matching that integrates different strategies and LLMs. In future work, we will further uncover the feasibility of automated composition and extend this paradigm to different stages of entity resolution.
Ethical Considerations
We used publicly available LLMs and datasets to conduct experiments in this paper. Our code implementation and application scenarios do not involve any ethical issues.
@@
Appendix A Detailed Results of Open-Source LLMs on Different Strategies
Table 4, Table 5, Table 6, Table 7, Table 8, Table 9, Table 10, Table 11, and Table 12 show the performance of the experimented LLMs with three strategies on all datasets.
Appendix B Ablation Study on Each Dataset
@@
Figure 9 shows the performance of ComEM with respect to the different number of candidates retained for further selection. Similar to Figure 8, recall increases and precision decreases as $k$ increases. For some datasets such as Walmart-Amazon, DBLP-ACM, and TMDB-TVDB, F1 achieves the highest at $k=1$. For some of the other datasets, F1 changes dramatically as $k$ goes from 2 to 5. How to choose $k$ to balance the LLM capabilities and the actual situation is a direction that could be explored. One possible solution might be to dynamically adjust the number according to the probabilities of comparisons for each record.

### v2

@@ 5 Conclusion
In this paper, we investigate three strategies for entity matching using LLMs to bridge the gap between local matching and global consistency of ER. Our research shows that incorporating record interactions is essential for LLM-based entity matching. By examining the effect of broad LLMs on these strategies, we further design a ComEM framework that integrates the advantages of multiple strategies and LLMs. The effectiveness and cost efficiency of ComEM highlight the importance of task decomposition and LLM composition, opening up new avenues for entity matching using LLMs.
@@
This study aims to investigate different strategies for LLM-based entity matching. We conducted thorough experiments with 1 commercial LLM and 8 open-source LLMs to provide a broad base for our analysis. The selection of models is based on considerations of popularity, availability, and cost. Future research could explore whether similar findings hold as LLMs evolve and how performance changes relative to our results.
Since LLMs have been trained on massive web data, they are likely to have seen similar and same records, or even some matching results, even though the labels of the matches are stored separately. Nevertheless, the performance of these strategies is relatively consistent across 9 LLMs and varies greatly for the same LLM when using different strategies, highlighting that data exposure is not the determining factor in their effectiveness. In the future, it will be valuable to evaluate LLM-based entity matching on new or non-public data.
The investigation of different strategies was conducted using basic zero/few-shot prompting, a simple and effective paradigm for applying LLMs. We could not ignore the role of potential advanced prompt engineering methods in improving the accuracy and robustness of LLMs. In addition, fine-tuning LLMs for better execution of different strategies is also a worthwhile direction.
Finally, we have demonstrated the effectiveness of the compound framework in entity matching that integrates different strategies and LLMs. We would like to continue to develop specific modules for entity matching and extend this paradigm to different stages of entity resolution.
Appendix A Prompts
The prompts for various strategies of LLM-based entity matching used in this paper are presented in Table 4. To ensure fairness, the same prompts were used for all experimental LLMs. These prompts were constructed through a manual process of prompt engineering, which involved the testing and comparing of multiple variations to determine the most effective ones. In addition to the task description, we included specific response instructions such as “Answer "Yes" if they do and "No" if they do not” to guide the responses of LLMs. For in-context learning, prompts and labels were repeatedly inputted for each example, followed by the records to be matched. We post-processed the LLM responses to obtain the final predicted labels.
Matching
Do the two entity records refer to the same real-world entity? Answer "Yes" if they do and "No" if they do not.
Record 1: {{ record_left }}
Record 2: {{ record_right }}
Comparing
Which of the following two records is more likely to refer to the same real-world entity as the given record? Answer with the corresponding record identifier "Record A" or "Record B".
Given entity record: {{ anchor }}
Record A: {{ candidate_left }}
Record B: {{ candidate_right }}
Selecting
Select a record from the following candidates that refers to the same real-world entity as the given record. Answer with the corresponding record number surrounded by "[]" or "[0]" if there is none.
Given entity record: {{ anchor }}
Candidate records:{for candidate in candidates %}
[{{ loop.index }}] {{ candidate }}{endfor %}
Table 4: Specific prompts of different strategies. We use JinJa template syntax to display the placeholders for the anchor record and potential matches (candidates).
@@ Appendix A Prompts
Appendix B Detailed Results of Open-Source LLMs under Different Strategies
We provide the detailed F1 scores of open-source LLMs under different strategies in Table 5. Among the 8 LLMs evaluated in our experiment, 6 achieve the best performance through the selecting strategy, and 2 achieve better performance through the comparing strategy. In summary, our proposed strategies are universally applicable across different LLMs for entity matching. We have observed that it is difficult to limit the output of many chat-tuned LLMs simply by prompts, which may affect their actual performance in entity matching. Therefore, how to calibrate the label probabilities from the long-form responses of LLMs is also important for performance improvement.
LLM Strategy AB AG DA DS IM IV TT WA Mean
Mistral-Instruct-7B Matching 40.70 37.77 24.68 28.89 64.86 64.49 49.91 55.96 45.91
Comparing 54.68 32.10 49.28 49.75 74.38 52.25 81.69 44.39 54.82
Selecting 67.26 57.31 83.36 74.27 87.84 76.95 80.89 62.54 73.80
Mistral-Instruct-8x7B Matching 77.67 34.76 67.20 60.09 82.26 53.57 72.99 50.57 62.39
Comparing 67.81 25.20 81.48 75.54 75.15 54.05 73.93 41.22 61.80
Selecting 79.58 61.16 85.05 79.37 90.34 77.15 81.23 78.84 79.09
Solar-Instruct-10.7B Matching 68.80 45.60 47.02 38.32 70.35 40.49 75.18 70.57 57.04
Comparing 86.22 49.14 84.70 75.16 61.68 32.57 77.49 74.41 67.67
Selecting 74.27 62.05 74.93 65.50 79.56 59.68 73.96 74.89 70.60
Flan-T5-XXL (11B) Matching 77.85 58.35 87.63 80.34 71.82 51.62 74.62 67.23 71.18
Comparing 84.21 56.85 94.49 85.82 65.33 49.88 84.28 67.89 73.60
Selecting 77.52 69.83 84.77 80.29 85.07 68.05 78.90 77.33 77.72
Flan-UL2 (20B) Matching 83.39 52.73 81.97 67.53 82.35 40.56 70.88 74.07 69.19
Comparing 88.09 64.52 94.81 88.26 71.43 39.51 83.66 80.66 76.37
Selecting 80.34 71.82 84.00 80.57 84.09 65.70 80.99 71.94 77.43
Command-R-35B Matching 49.87 32.87 47.87 44.46 91.45 69.69 63.14 36.81 54.52
Comparing 72.31 51.27 76.82 65.91 90.91 77.00 86.09 57.24 72.20
Selecting 78.16 65.52 83.67 79.54 85.26 75.33 79.06 80.58 78.39
Llama-3-8B Matching 31.01 21.97 19.27 19.27 44.78 40.24 31.55 23.91 29.00
Comparing 80.06 61.27 84.85 72.54 80.13 76.36 79.82 80.29 76.91
Selecting 74.37 49.50 78.91 68.79 76.27 54.77 69.66 42.33 64.33
Qwen-2-7B Matching 63.41 47.33 68.35 52.46 82.89 55.54 71.84 55.06 62.11
Comparing 84.32 56.88 88.78 76.57 93.17 65.07 86.50 75.39 78.34
Selecting 72.39 61.03 81.49 76.57 82.97 73.48 78.55 72.96 74.93
Table 5: F1 score of open-source LLMs under different strategies.
Appendix C Ablation Study on Each Dataset
@@
Figure 8 shows the performance of ComEM with respect to the varying number of top candidates retained for further selection. Similar to Figure 7, recall increases and precision decreases as $k$ increases. For the simplest dataset “DBLP-ACM”, F1 achieves the highest value at $k=1$. For some other datasets, F1 changes dramatically as $k$ goes from 2 to 5. How to tune $k$ to balance the LLM capabilities and the actual situation is a direction that could be explored. One possible solution might be to dynamically adjust the number according to the similarity scores of potential matches.